Suppose you are standing on a boat in the ocean looking out at a person swimming in the water. The waves swell under your boat at a frequency of 0.80 Hz, while traveling with a wave speed of 4.5 meters per second. At any given moment, you count that there are exactly 6 complete waves between you and the swimmer. How far away is the swimmer? answer in meters.
1
Expert's answer
2011-04-06T13:40:35-0400
We shoud find the wavelength of the waves:
λ = v/f = 4.5 / 0.8 = 5.625 m.
6 λ = 6* 5.625 = 33.75 m.
The distance is& 33.75 m.